An annual corporate bond is priced to yield 7.5% annually and has a price of $972.18. Its Macaulay duration is 5.8215. If the rate decreases to 7.4%, estimate the price of the bond using
(a) first-order modified approximation;
(b) first-order Macaulay approximation.
Answer #1

Solution :

It is given that the Macaulay duration = 5.8215

Modified duration = Macaulay duration / ( 1 + yield ) = 5.8215/(1+7.5%) = 5.8215/1.075 = 5.4153

It is given that the change in yield = 0.1%

Part A ) change in price using modified duration

Change in price /price = - duration *change in yield = -5.4153 * -0.001

Change in price = 972.18 * 5.4153* 0.001 = 5.26

Price will increase by 5.26 and new price = 972.18+ 5.26 = 977.44

Part B )

Change in price using Macaulay duration

Change in price / price = - duration * change in yield = -5.8215* -0.001

Change in price = 972.18 * 5.8215 * 0.001 = 5.66

Price will increase by 5.66 and new price will be 972.18+5.66 = 977.84